Find the derivative of y= 4tan5α

1
Expert's answer
2021-10-13T10:17:58-0400

Given

"Y=4tan5\alpha............(1)"

We know that

"\frac{d}{da}(tanm\alpha)=sec^2(m\alpha).\frac{d}{d\alpha}(ma)\\=sec^2(m\alpha)=(mx_1)"

"\implies\frac{d}{d\alpha}(tanm\alpha)=msec^2(m\alpha)..................(2)"


Now differentiating (1) wrt "\alpha" both sides we get

"\frac{dY}{d\alpha}=\frac{d}{d\alpha}(4tan5\alpha)=4.\frac{d}{d\alpha}(tan5\alpha)\\=4\times5sec^2(5\alpha): using(2)"


"\frac{dY}{d\alpha}=20sec^2(5\alpha)"